# FoREST: Frame of Reference Evaluation in Spatial Reasoning Tasks

### v1

@@ 5.3 Results
Figure 3: Confusion matrices of spatial relation answers when Llama3 and GPT-4o are required to adapt FoR in the 0-shot and (SG+CoT) settings.
@@
RQ2. Can the model adapt FoR when answering the questions? To address this research question, we analyze QA that required FoR comprehension results in C-Split from Table 2. Note that the context and question in these tasks explicitly indicate a perspective. The results indicate that LLMs struggle with FoR conversion, particularly when the question has relatum and the context has camera perspectives, achieving only up to 55.5% accuracy. We further demonstrate how Llama3 and GPT-4o adapt FoR using the confusion matrix in Figure 3. Our findings reveal that pure-text LLM (Llama3) has confusion between left and right. Humans typically reverse front and back while preserving left and right when describing the spatial relation from perspective. However, Llama3 incorrectly reverses left and right, leading to poor adaptation to the camera perspective. In contrast, very large multimodal-language models like GPT-4o follow the expected pattern, as observed by Zhang et al. 2025. While our GPT-4o results suggest some ability to convert the relatum’s perspective into the camera’s with in-context learning (72% accuracy), the reverse transformation in the textual domain remains challenging (53% accuracy). This difficulty persists when converting spatial relations from the camera perspective from images to the relatum’s perspective as observed in Zhang et al. 2025.
@@
RQ3. How can an explicit FoR identification help spatial reasoning in QA? We compare CoT and CoT+SG results to evaluate how explicit FoR identification affects LLMs’ spatial reasoning in QA. Based on C-Split results (Table 3), incorporating SG encourages the model to identify the perspective from input expression ranging from 2.9% to 30% in cases where the context and question share the same perspective. These cases are easier as the models do not need FoR adaptation. With one exception, as can be seen, the performance level of the Llama3 baseline, for questions with camera perspective, is so poor that FoR identification in SG cannot help boost its performance compared to the improvements made in other language models. We should note that among our selected LLMs, Llama3 is the only one not trained with visual information; we speculate this can be a factor in LLMs’ understanding of FoR. SG is less effective in reasoning when the context and question have different perspectives despite aiding in identifying the correct FoR of the spatial description in the context. This limitation is evident in A-Split results (Table 1), where models only show significant improvement when SG aligns their preference with the question’s perspective, as seen in Qwen2-72B and GPT-4o. SG identification results are reported in the Appendix E. Still, incorporating FoR identification improves overall spatial reasoning (see the Avg column for SG+CoT in Table 3).
@@
RQ4. How can explicit FoR identification help spatial reasoning in visualization? We evaluate SG layout diffusion to assess the impact of incorporating FoR in image generation. We focus on VISORcond metric, as it better reflects the model’s spatial understanding than the overall performance measured by VISORuncond, which is reported in Appendix C due to space limitation. Table 3 shows that adding spatial information and FoR classes (SG+GLIGEN) improves performance across all splits compared to the baseline models (GLIGEN). SG improved the model’s performance when expressions can be interpreted as relative FoR. These results align with the QA results shown in Table 1 indicating that Llama3 prefers relative FoR if dealing with the camera’s perspective. In contrast, baseline diffusion models (SD-1.5 and SD-2.1) perform better for intrinsic FoR even though GLIGEN is based on SD-2.1. This outcome might be due to GLIGEN’s reliance on bounding boxes for generating spatial configurations, which makes it struggle with intrinsic FoR due to the absence of object properties and orientation. Nevertheless, incorporating FoR information via SG-prompting improves performance across all FoR classes despite this specific bias. We provide further analysis on SG for the layout generation in Appendix D.
@@ 6 Related Work
Frame of Reference in Cognitive Studies. The concept of the frame of reference in cognitive studies was introduced by Levinson 2003 and later expanded with more diverse spatial relations (Tenbrink, 2011). Subsequent research investigated the human preferences for specific FoR classes (Edmonds-Wathen, 2012; Vukovic and Williams, 2015; Shusterman and Li, 2016; Ruotolo et al., 2016). For instance, Ruotolo et al. 2016 examined how FoR influences scene memorization and description under time constraints. Their study found that participants performed better when spatial relations were based on their position rather than external objects, highlighting a distinction between relative and intrinsic FoR.
Frame of Reference in AI. Several benchmarks have been developed to evaluate the spatial understanding of AI models in multiple modalities; for instance, textual QA (Shi et al., 2022; Mirzaee and Kordjamshidi, 2022; Rizvi et al., 2024), and text-to-image (T2I) benchmarks (Gokhale et al., 2023; Huang et al., 2023; Cho et al., 2023a, b). However, most of these benchmarks overlook the frame of reference (FoR), assuming a single FoR for all instances despite its significance in cognitive studies. Recent works in vision-language research are beginning to address this problem. For instance, Liu et al. 2023 examines FoR’s impact on visual question-answering but focuses only on limited FoR categories. Our work covers more diverse FoRs. Zhang et al. 2025 examine FoR ambiguity and understanding in vision-language models by evaluating spatial relations derived from visual input under different FoR questions. Their approach relies on images from camera perspectives, with FoR indicated in the question. In contrast, our work focuses on the reasoning of spatial relations when dealing with multiple FoRs and when there are changes in perspective in explaining the context beyond the camera’s viewpoint. Additionally, we show that explicitly identifying the FoR helps improve spatial reasoning in both question-answering and text-to-image generation, particularly when involving multiple perspectives.
@@ 7 Conclusion
Given the significance of spatial reasoning in AI applications, we introduce Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark to evaluate Frame of Reference comprehension in textual spatial expressions via question-answering and grounding in visual modality by diffusion models. Based on this benchmark, our results reveal notable differences in FoR comprehension across LLMs and their struggle with questions that require adaptation between multiple FoRs. Moreover, the bias in FoR interpretations impacts the layout generation with LLMs for text-to-image models. To improve FoR comprehension, we propose Spatial-Guided prompting, which first generates spatial relation’s topological, distal, and directional type information in addition to FoR and includes this information in downstream task prompting. Employing SG improves the overall performance in QA tasks requiring FoR understanding and text-to-image generation.
@@ 8 Limitations
While we analyze LLMs’ shortcomings, our benchmark only highlights areas for improvement, not harming the model. The trustworthiness and reliability of the LLMs are still a research challenge. Our analysis is confined to the spatial reasoning domain and does not account for biases related to gender or race. However, we acknowledge that linguistic and cultural variations in spatial expression are not considered, as our study focuses solely on English. Extending this work to multiple languages could reveal important differences in FoR adaptation. Our analysis is still limited to the synthetic environment. Future research should consider the broader implications of the frame of reference of spatial reasoning in real-world applications. Additionally, our experiments require substantial GPU resources, limiting the selection of LLMs and constraining the feasibility of testing larger models. The computational demands also pose accessibility challenges for researchers with limited resources. We find no ethical concerns in our methodology or results, as our study does not involve human subjects or sensitive data.
Appendix A Dataset Statistics
The FoREST dataset statistic is provided in the Table 4.
Case A-Split A-Split for T2I FoR class C-Spilt C-split for T2I
Cow Case 792 3168 External Relative 1528 4288
Box Case 120 120 External Intrinsic 920 3680
Car Case 128 512 Internal Intrinsic 128 0
Pen Case 488 488 Internal Relative 248 0
Total 1528 4288 Total 2824 7968
Table 4: Dataset Statistic of FoREST dataset.
Appendix B Details Creation of FoREST dataset
@@
Based on the selected locatum and relatum. To create an A-split spatial expression, we substitute the actual locatum and relatum objects in the Spatial Relation template. After obtaining the A-split contexts, we create their counterparts using the perspective/topology clauses to make the counterparts in C-spilt. Then, we obtain the I-A and I-C split by applying the directional template to the first occurrence of relatum when it has intrinsic directions. The directional templates are "that is facing towards," "that is facing backward," "that is facing to the left," and "that is facing to the right." All the templates are in the Table 6. We then construct the scene configuration from each modified spatial expression and send it to the simulator developed using Unity3D. Eventually, the simulator produces four visualization images for each scene configuration.
Category Object Intrinsic Direction Container
@@
bog object with intrinsic directions bench, chair
@@ B.1 Simulation Details
B.2 Object Models and Background
@@
Figure 4: All 3d models used to generate visualizations for FoREST.
@@
B.3 Textual templates
All the templates used to create FoREST are given in Table 6.
Appendix C VISOR uncond Score
@@
VISORuncond provides the overall spatial relation score, including images with object generation errors. Since it is less focused on evaluating spatial interpretation than VISORcond, which assesses explicitly the text-to-image model’s spatial reasoning, we report VISORuncond results here in the Table 7 rather than in the main paper. The results are similar to the pattern observed in VISORuncond that the based models(SD-1.5 and SD-2.1) perform better in the relative frame of reference, while the layout-to-image models, i.e., GLIGEN, are better in the intrinsic frame of reference.
@@
Table 7: VISORuncond score on the A-Split and C-Split where $I$ refer to the Cow Case and Car Case where relatum has intrinsic directions, and $R$ refer to the Box Case and Pen case where relatum lacks intrinsic directions, $avg$ is mirco-average of $I$ and $R$. cond and uncond are explained in Section 5.1.
Appendix D Analyze the improvements in SG-prompting for T2I.
@@
Appendix E Frame of Reference Identification
We evaluate the LLMs’ performance in recognizing the FoR classes from given spatial expressions. The LLMs receive spatial expression, denoted as $T$, and output one FoR class, $FoR$, from the valid set of FoR classes, $For\in$external relative, external intrinsic, internal intrinsic, internal relative}. All in-context learning examples are in the Appendix F.
E.1 Experimental Setting

### v2

@@ 5.3 Results
Figure 3: Confusion matrices of predicted spatial relations by Llama3 and GPT-4o when adapting to FoR is required. 0-shot and SG-COT settings are reported here.
@@
RQ2. Can LLMs adapt FoR when answering the questions? To address this RQ, we analyze QA results of C-split in Table 2, where context and question explicitly specify perspective. Results show that LLMs struggle with FoR conversion, achieving only up to 55.5% accuracy when the question uses the relatum’s perspective and the context uses the camera’s. We further examine how Llama3 and GPT-4o adapt FoR using the confusion matrix in Figure 3. Our findings reveal that pure-text LLM (Llama3) reverses left and right. This contrasts with humans when describing the spatial relation from a perspective in English, which reverses front and back while preserving lateral directions. This leads to poor adaptation to the camera perspective of Llama3. In contrast, large multimodal models like GPT-4o follow expected patterns, consistent with Zhang et al. 2025. While our GPT-4o results suggest some abilities to convert the relatum’s perspective into the camera’s with in-context learning (72% accuracy), the reverse transformation in the textual domain remains challenging (53% accuracy). A similar trend appears when comparing Qwen2 and Qwen2VL, as discussed in Appendix B. Qwen2VL performs better when shifting to the camera perspective but performs worse with reverse transformation, often failing to generate correct reasoning. This difficulty persists for several models when converting spatial relations from images to relatum’s perspective, as noted in Zhang et al. 2025. A possible reason is that the model often overlooks orientation tokens, critical for answering FoR questions, based on Attention analysis provided in Appendix C.
@@
RQ3. How can an explicit FoR identification help spatial reasoning in QA? We compare CoT and CoT+SG results to assess the effect of FoR identification on LLMs’ spatial reasoning in QA. Based on C-Split results (Table 2), incorporating SG improves the model’s ability to identify the correct perspective from input expression ranging from 2.9% to 30% of cases where the context and question share the same perspective. These cases are easier as the models do not need FoR adaptation. Two notable exceptions emerge. First, Llama3 performs poorly on camera-perspective questions, and FoR identification via SG fails to improve its performance. This may be due to Llama3’s lack of visual training, which we speculate limits its FoR understanding. Second, Qwen2VL struggles with relatum-perspective reasoning, showing negative gains even with CoT. SG is less effective when context and question differ in perspective; while it helps identify the correct FoR in context, it does not enhance reasoning across perspectives. This limitation is evident in A-Split results (Table 1), where models only improve significantly when SG aligns their preference with the question’s perspective, as seen in Qwen2-72B and GPT-4o. SG identification results are reported in the Appendix H. Still, FoR identification improves overall spatial reasoning (see Avg. column for SG+CoT in Table 3).
@@
RQ4. How can explicit FoR identification help spatial reasoning in visualization? We evaluate SG layout diffusion to assess the impact of incorporating FoR in image generation. We focus on VISORcond metric, which better reflects the model’s spatial understanding than the overall performance measured by VISORuncond, reported in Appendix F. Table 3 shows that adding spatial information and FoR classes (SG+GLIGEN) improves performance across all splits compared to the baseline models (GLIGEN). In particular, SG improved the model’s performance when expressions follow a relative FoR. This finding aligns with QA results in Table 1, where Llama3 prefers relative FoR in camera-perspective scenarios. In contrast, baseline diffusion models (SD-1.5 and SD-2.1) perform better for intrinsic FoR, even though GLIGEN is based on SD-2.1. This outcome may be due to GLIGEN’s reliance on bounding boxes for spatial configurations, which makes it struggle with intrinsic FoR due to the absence of object properties and orientation. Despite this bias, incorporating FoR information via SG-prompting improves performance across all FoR classes. We provide further analysis on SG for the layout generation in Appendix G.
@@ 6 Related Work
Frame of Reference in Cognitive Studies. The concept of the frame of reference in cognitive studies was introduced by Levinson 2003 and later expanded with more diverse spatial relations (Tenbrink, 2011). Subsequent research investigated the human preferences for specific FoR classes (Vukovic and Williams, 2015; Shusterman and Li, 2016; Ruotolo et al., 2016; Coventry et al., 2018). For instance, Ruotolo et al. 2016 examined how FoR influences scene memorization. Their study found that participants performed better when spatial relations were based on their position rather than external objects, highlighting a distinction between relative and intrinsic FoR.
Frame of Reference in AI. Several benchmarks have been developed to evaluate the spatial understanding of AI models in multiple modalities; for instance, textual QA (Shi et al., 2022; Mirzaee and Kordjamshidi, 2022; Rizvi et al., 2024), and text-to-image (T2I) benchmarks (Gokhale et al., 2023; Cho et al., 2023a, b). However, most of these benchmarks overlook FoR, assuming a single FoR for all instances despite its significance in cognitive studies. Recent vision-language studies have begun addressing this gap Liu et al. (2023); Zhang et al. (2025). For instance, Zhang et al. 2025 explores FoR ambiguity by evaluating spatial relations from camera-perspective images, with FoR specified in the question. In contrast, our work focuses on the spatial reasoning when dealing with multiple FoRs and when there are changes in perspective in explaining the context beyond the camera’s viewpoint. We further show that explicitly identifying FoR enhances spatial reasoning in QA and T2I.
@@ 7 Conclusion
Given the significance of spatial reasoning in AI, we introduce Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark to evaluate Frame of Reference comprehension in textual spatial expressions through question-answering and grounding in visual modality via diffusion models. Using this benchmark, we identify notable differences in FoR comprehension across LLMs and their struggle with questions that require adaptation between multiple FoRs. Moreover, the bias in FoR interpretations impacts the layout generation with LLMs for text-to-image models. To improve FoR comprehension, we propose Spatial-Guided prompting, which extracts spatial relations’ topological, distal, and directional type information in addition to FoR and includes this information in downstream task prompting. Employing SG improves performance in both QA tasks requiring FoR understanding and text-to-image generation.
@@ 8 Limitations
While we analyze LLMs’ shortcomings, our benchmark only highlights areas for improvement. The trustworthiness and reliability of the LLMs are still a research challenge. Our analysis is confined to the spatial reasoning domain and does not account for biases related to gender or race. However, we acknowledge that linguistic and cultural variations in spatial expression are not considered, as our study focuses solely on English. Extending this work to multiple languages could reveal important differences in FoR adaptation. Our analysis is still limited to the synthetic environment. Future research should consider the broader implications of the frame of reference of spatial reasoning in real-world applications. Additionally, our experiments require substantial GPU resources, limiting the selection of LLMs and constraining the feasibility of testing larger models. The computational demands also pose accessibility challenges for researchers with limited resources. We find no ethical concerns in our methodology or results, as our study does not involve human subjects or sensitive data.
Appendix A Additional details of FoREST Dataset
Figure 4: All 3d models used to generate visualizations for FoREST.
A.1 Dataset Statistics
The FoREST dataset statistic is provided in the Table 4 for A-split and Table 5 for C-split.
A.2 Details Creation of FoREST dataset
@@
Based on the selected locatum and relatum. To create an A-split spatial expression, we substitute the actual locatum and relatum objects in the Spatial Relation template. After obtaining the A-split contexts, we create their counterparts using the perspective/topology clauses to make the counterparts in the C-split. Then, we obtain the I-A and I-C split by applying the directional template to the first occurrence of relatum when it has intrinsic directions. The directional templates are "that is facing towards," "that is facing backward," "that is facing to the left," and "that is facing to the right." All the templates are in the Table 7. We then construct the scene configuration from each modified spatial expression and send it to the simulator developed using Unity3D. Eventually, the simulator produces four visualization images for each scene configuration.
Category Object(s) Intrinsic Direction Possible Container
@@
Big object with intrinsic directions bench, chair
@@ A.2.1 Simulation Details
A.2.2 Object Models and Background
@@
Table 7: All templates used to create FoREST dataset.
A.2.3 Textual templates
All manually created templates used to create FoREST are given in Table 7.
@@ A.2.3 Textual templates
A.2.4 GPT template
We prompt GPT with an initial question using a specific template (the first template in Table 8) to generate four additional variations. Each generated template was manually reviewed and validated before being added to the set of questions in our corpus. All templates are listed in Table 8. Question templates for asking the spatial relation from relatum view are edited by replacing references from the camera to the relatum object.
Appendix B Comparison between Qwen2 and Qwen2-VL in QA setting
Figure 5: Confusion matrices of spatial relation answers when Qwen2 and Qwen2-VL must adapt FoR in the 0-shot and (SG+CoT) settings.
According to the A-split results in Table 1, Qwen2VL demonstrates a more balanced preference pattern and reasonable performance compared to Qwen2. This suggests that, unlike Qwen2, which often extracts spatial relations without accounting for perspective, Qwen2VL appears capable of reasoning about perspective shifts as well. This observation is further supported by the C-split results in Table 2, where Qwen2VL significantly outperforms Qwen2 in cases requiring adapting the spatial relations mentioned between different perspectives in most settings. These results support our hypothesis that visual training enhances the model’s ability to interpret perspective changes.
However, despite these improvements, Qwen2VL still struggles to provide reasoning from the relatum’s perspective on questions in CoT and SG+CoT settings. This suggests that while the model leverages visual training to reason about perspective and perform perspective shifts when necessary, its reasoning is more accurate when questions are framed from the camera’s viewpoint. However, it struggles with questions asked from the relatum’s perspective. This limitation is expected, as training data—particularly image captioning datasets—are typically annotated from a human, camera-based perspective, restricting the model’s ability to generalize across different viewpoints.
Appendix C Attention Analysis
Figure 6: Attention weights of input context for Question Answering in A-split. Considering the following semantic concepts: locatum mention in context, relation in context, relatum in context, relatum orientation, perspective in question, locatum in question, and relatum in question. Three patterns are considered: (1) incorrect answers, (2) correct answers assuming a shared perspective between context and question, and (3) correct answers assuming different perspectives.
C.1 Experiment Setting
We conducted an attention-based interpretability analysis using Qwen2 in a 0-shot setting, where only the context and question were provided. We focused on analyzing attention weights across key semantic elements in both A-split and C-split. We analyzed the attention weights for various groups of tokens that convey the following concepts: locatum mention occurring in the context, relation in context, relatum in context, perspective in context, relatum orientation, perspective in question, locatum in question, and relatum occurring in the question.
We categorized response patterns into three scenarios: (1) the model answers incorrectly, (2) the model answers correctly with the assumption that the same perspective holds for context and question, and (3) the model answers correctly with the assumption that context and question have different perspectives. Note that we separate the analysis for both splits because the conceptual tokens are different, that is, the perspective mentioned in the context is only for the C-split.
C.2 Results
A-split. According to Figure 6, we found that, for the correct responses, the model paid the least attention to the question’s perspective in the same perspective. Meanwhile, for the different perspective case, attention significantly increased to tokens related to the relation, the question perspective, and the relatum in the question. This result illustrates that these tokens played an important role in adapting spatial relations for perspective shifts. However, when the model increased attention weight on relatum orientation compared to other tokens, more errors occurred in answering the questions.
(a) Questions that require changing perspective.
(b) Questions that do not require changing perspective.
Figure 7: Attention weights of input context for Question Answering in C-split. Considering the following semantic concepts: locatum mention in context, relation in context, relatum in context, perspective in context, relatum orientation, perspective in question, locatum in question, and relatum in question. Two patterns are considered: (1) incorrect answers, (2) correct answers. We separate the questions that require and not-require perspective changing.
C-split. According to Figure 7, in cases that require a shift in perspective, we observed that the model pays significantly more attention to the spatial relation in context, the question perspective, and the relatum in the question concepts necessary for answering correctly. However, the model largely overlooks the orientation tokens, which also play a crucial role in resolving these questions. This lack of attention to orientation may majorly contribute to the model’s failures in perspective-shifting scenarios. In cases where the relation mentioned remains the same in both the context and the answer, we observed that the model, in incorrect predictions, tends to assign higher attention to the relatum in the question. In contrast, correct predictions are associated with more balanced attention across tokens. Maintaining moderate attention to all tokens appears to help the model consistently respond with the correct relation, repeating it from the context.
Appendix D Quantitative Examples
D.1 Example of QA Bias in the A-split
To illustrate example biases in the A-split, we present three types of model behavior. The first, shown in Figure 8, occurs when the model assumes a shared perspective between context and question, allowing it to extract spatial relations directly without FoR reasoning. This scenario is the most common behavior of Qwen2. The second, illustrated in Figure 9, arises when the model assumes the context always reflects the relatum’s perspective, favoring intrinsic over relative interpretation. Lastly, Figure 10 shows the model assuming the context always adopts the camera perspective.
Figure 8: The first example in A-split of FoREST, where the model correctly answers both perspectives. In this example, the model responds with the spatial relation in the context, assuming all questions and context have the same perspective.
Figure 9: The second example in A-split of FoREST, where the model correctly answers both perspectives. In this example, the model assumes the context has a relatum perspective.
Figure 10: The third example in A-split of FoREST, where the model correctly answers both perspectives. In this example, the model assumes the context has a camera perspective.
D.2 Incorrect QA reasoning in the C-split
Our results demonstrate that models struggle with converting between different perspectives when required to answer the questions, especially when the correct answer is either left or right. This difficulty contributes to their performance, as discussed in research question 2 in the paper (RQ2). To further illustrate LLMs’ struggle, we provide quantitative examples of failure cases, focusing specifically on the models’ most challenging instances in Figure 11.
Figure 11: An example of an incorrect SG+CoT case, where the model generates incorrect reasoning and mistakenly converts left to right.
Appendix E Template Variation in Question-Answering
We report average accuracy on the C-split across all question templates for Qwen2 and GPT-4o in Table 9. While results show notable variation across templates, no single template consistently outperforms others across all tasks. Template effectiveness depends on both the prompting technique and the underlying model. Simpler templates generally yield higher accuracy, though more complex expressions may reduce overall performance. Nevertheless, the comparative trend remains consistent, as SG+CoT often outperforms CoT in overall improvement across both models. Notably, SG prompting yields the smallest performance variation across templates.
Model T0 T1 T2 T3 T4 Avg. Highest $\Delta$
Qwen2 (0-shot) 70.23 61.36 69.35 65.55 67.97 66.91 8.87
Qwen2 (4-shot) 65.15 64.01 68.80 63.24 65.46 65.34 5.57
Qwen2 (CoT) 71.26 66.11 72.78 63.40 71.87 69.10 9.37
Qwen2 (SG + CoT) 71.15 70.92 72.89 70.11 71.69 71.36 2.78
GPT-4o (0-shot) 64.02 61.14 73.16 65.77 56.66 64.19 16.50
GPT-4o (4-shot) 65.85 52.23 70.55 61.64 58.33 61.78 18.31
GPT-4o (CoT) 80.98 69.65 84.81 83.24 74.98 78.78 15.16
GPT-4o (SG + CoT) 85.74 85.02 84.10 86.53 87.57 85.79 3.47
Table 9: Average accuracy across different templates. Ti refers to the generated QA template $i$ in Table 7. $\Delta$ measures the difference in accuracy between two different templates.
Appendix F VISOR-uncond Score
@@
VISORuncond provides the overall spatial relation score, including images with object generation errors. Since it is less focused on evaluating spatial interpretation than VISORcond, which assesses explicitly the text-to-image model’s spatial reasoning, we report VISORuncond results here in the Table 10 rather than in the main paper. The results are similar to the pattern observed in VISORuncond that the based models(SD-1.5 and SD-2.1) perform better in the relative frame of reference, while the layout-to image models, i.e. GLIGEN, are better in the intrinsic frame of reference.
@@
Table 10: VISORuncond score on the I-A-Split and I-C-Split where $I$ refer to the Cow Case and Car Case where relatum has intrinsic directions, and $R$ refer to the Box Case and Pen case where relatum lacks intrinsic directions, $avg$ is mirco-average of $I$ and $R$. cond and uncond are explained in Section 5.1.
Appendix G Analyze the improvements in SG-prompting for T2I.
@@
Appendix H Frame of Reference Identification
We evaluate the LLMs’ performance in recognizing the FoR classes from given spatial expressions. The LLMs receive spatial expression, denoted as $T$, and output one FoR class, $FoR$, from the valid set of FoR classes, $For\in$external relative, external intrinsic, internal intrinsic, internal relative}. All in-context learning examples are in the Appendix K.
H.1 Experimental Setting